# AICO: Feature Significance Tests for Supervised Learning

### v3

@@ 2.2 Multiple features, panel data and missing data
We have focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The approach naturally extends to testing multiple features, i.e. subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions.
Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains. Appendix B discusses two simple extensions—conditioning and grouping—to construct the feature effect for panel data.
Real-world data sets often have missing feature values. Appendix C shows how to treat missing or imputed values when constructing $\Delta_{i}$.
@@ 3 Hypothesis testing
To construct a hypothesis test, we consider the empirical distribution of the feature effect $\Delta_{i}$ across the test samples $i\in I_{2}$ (the feature $\ell\in[d]$ to be tested remains fixed). If that distribution has a non-positive center, then the feature is deemed insignificant. A large positive center is evidence of a feature being significant. There is a battery of alternative tests for the mean, median and other statistics, including $t$, sign, Wicoxon, and permutation tests. We propose a simple sign test, which requires minimal assumptions on the data yet offers meaningful asymptotic relative efficiency vs. a $t$-test.
@@
The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$ with unique median $M\in\mathbb{R}$. The associated distribution function $F_{\Delta}$ is continuous. [^7] A sufficient condition for the uniqueness of $M$ is strict monotonicity of $F_{\Delta}$ in a neighborhood of $M$.
Let $B_{n,p}$ be a Binomial-$(n,p)$ variable and let $q_{\lambda}(B_{n,p})=\inf\{x\geq 0:P(B_{n,p}\leq x)\geq\lambda\}$ be its (lower) quantile at $\lambda\in(0,1)$. Let $N=|I_{2}|$ be the test sample size. Moreover, let $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ be the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
@@
For brevity, we often suppress the dependence of the test statistic $n_{+}$ on $M_{0}$. The test rejects $H_{0}$ with probability one if $n_{+}$, the number of feature effects in the test set larger than $M_{0}$, exceeds the $(1-\alpha)$-quantile of the Binomial-$(N,1/2)$ distribution. The test randomizes in case $n_{+}$ hits the boundary of the rejection region. Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, randomization renders the test more aggressive; it ensures maximum power and guarantees exactness in the sense that
@@
Algorithm 1 details the implementation of the test. Our applications of the test in Sections 5 and 6 below will use the value $M_{0}=0$ in (1).
@@ 4 Measuring feature importance
We require measures of feature importance that complement the feature $p$-values delivered by our test. A local measure of feature importance is given by the feature effect $\Delta_{i}$, which represents the impact of a feature at the sample level. The median $M$ of the (conditional) feature effect distribution $F_{\Delta}$ is a natural measure of feature importance at the population level. Under Assumption 3.1, a consistent estimator of $M$ is the empirical median
$\widehat{M}=\mbox{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$ and approximate equality holding for even $N$. Thus (6) is (approximately) as likely to overestimate $M$ as to underestimate it. Moreover, the distribution of $\widehat{M}$ is continuous. The estimator $\widehat{M}$ is unbiased if $F_{\Delta}$ is symmetric about $M$; see Section 4.4 of ?.
@@
We develop several useful properties of the population-level feature importance measure defined by the median $M$. To this end, let $Z=(X,Y)\sim\mathbb{P}$ where $\mathbb{P}$ denotes the population law from which the observed data points $Z_{i}$ are independently drawn. For concreteness, consider a model for estimating the conditional mean, and take $\mu(X)=\mathbb{E}(Y|X)$ and the score function $S$ as given. The (population) feature effect is given by $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$ where $\underline{X}$ and $\overline{X}$ are suitable baseline and comparison features, respectively.
As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers some degrees of freedom. Consider the basic choice from Example 2.1,
@@
where for convenience we assume, without loss of generality, feature $X^{1}$ is to be tested. First note that for the choice (10) and given $\mu$ and $S$, the feature effect $\Delta$ is completely determined by the marginal laws of the features and the joint law of the response $Y$ and $X^{1}$. Thus the feature dependence structure does not matter; the importance measure $M$ generated by (10) is robust to feature correlation. This is a remarkable property given the sensitivity of popular importance measures including SHAP and LOCO to correlation among features (e.g. ?). The correlation sensitivity induces a bias in these measures, which can lead to dubious importance rankings.
Next we explore the relationship between $M$ and conditional independence. ?, ?, ?, ?, ? and others develop tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?. We show that the hypothesis we test ($M\leq M_{0}=0$) is generally weaker than an alternative conditional independence hypothesis. Below, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
@@
$Y=\mu(X)+\epsilon,\;\text{where}\;\mu(X)=\beta\cdot X\;\text{and}\;\beta=(\beta^{1},...,\beta^{d})\in\mathbb{R}^{d}.$ (11)
Before stating a converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) (and symmetry of the distribution of $\Delta$), the median $M$ measuring feature importance is proportional to the square of the regression coefficient, which is the canonical measure of feature importance in a linear model. The proportionality factor is given by the variance of the feature.
@@ 5 Synthetic tasks
This section illustrates the robust performance of our test and feature importance measure on synthetic “known truth” regression and classification tasks using neural networks. It also benchmarks performance against existing model-agnostic tests and variable selection procedures, as well as alternative model-specific and -agnostic feature importance measures.
@@ 5.1 Regression
We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ of the form
@@
where the error $\epsilon$ is independent of $X$ and the conditional mean function $\mu$ is given by
@@
The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$. The features $X_{2},\ldots,X_{5},X_{7}$ are independent $N(0,1)$, $X_{8}$ is uniform on [1,1][-1,1], $X_{9}=1\{X_{2}+Z<0\}$ where $Z\sim N(0,1)$, $X_{10}$ is Poisson with mean 3 and $X_{11}$ and $X_{12}$ are independent variables with a $t$-distribution and 5 degrees of freedom. The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances and covariances equal to $0.85$, and $X_{17},X_{18}$ and $X_{19}$ are independent variables with a $t$-distribution and 5 degrees of freedom.
We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. We train a single-layer neural network model $f$ of the conditional mean $\mu(X)=\mathbb{E}(Y|X)$. Specifically,
@@
where $K$ is the number of units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated. The training procedure uses a standard $L_{2}$ loss criterion, $L_{2}$ regularization (of the weights $b_{k}$), the Adam optimization method, a batch size of 32, and 500 epochs with early stopping. We iterate across five random initializations. A validation set ($25\%$ of $I_{1}$, not used for training) and a Bayesian approach are used to optimize the regularization weight, the learning rate and the number of units, $K$. The optimal weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The $R^{2}$ on the training (and test) data exceeds $99\%$.
The testing procedure uses a score function $S(m,y)=-(m-y)^{2}$ consistent with the $L_{2}$ training criterion. We use the basic baseline and comparison feature specifications discussed in Example 2.1. That is, for testing feature $\ell\in[d]$ we take
$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i})$
$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$ (15)
where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous. For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$ and for the Poisson feature $\ell=10$, we use the “alternative mode” $v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1})$ where $p^{\ell}(\cdot|I_{1})$ is the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. The latter choice avoids baseline and comparison features being identical for too many samples, which is an issue arising with the standard mode. Our specifications yield baseline values in the support of $X^{9}$ and $X^{10}$.
@@
We run 10 independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.999Limited numerical precision can lead to ties, i.e. samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact is samples $\Delta_{i}$ with true, non-zero value being rounded to 0. We treat these samples as if they indeed had non-positive values not contributing to $n_{+}$, implying a conservative approach. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. The test has perfect power and does not produce false discoveries. Feature correlation (e.g. $\ell=1,6$) and discrete feature distributions ($\ell=9,10$) do not affect performance. The test also correctly detects the significance of features $\ell=2,4,5$ which influence the response only through product terms. Additional experiments show robustness to alternative baseline specifications (e.g. taking $v^{\ell}_{i}$ as the training mean also for the discrete features $\ell=9,10$, and using the median instead of the mean to construct $v^{\ell}_{i}$). We also verify robustness with respect to the score function $S$ including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
@@
Table 1 reports the results of the aforementioned alternative procedures for 10 independent training/testing experiments. CPT, HRT and LOCO deliver the same power as AICO but come with an increased false discovery rate (some but not all of these false discoveries can be addressed by a standard Bonferroni correction). The Gaussian Knockoff procedure selects the significant features 1 through 12 in each independent experiment. However, it does generate some false discoveries (which is consistent with control of the expected false discovery rate). The Deep Knockoff procedure reduces false discoveries at the expense of lower power. This approach has slight difficulty identifying the features entering the model (5.1) through the product term $2x_{4}x_{5}$.
We consider the performance of our AICO test as a function of the size $N=|I_{2}|$ of the test set. Table 2 reports the results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is reported in Table 1. The test offers maximum power even for $N$ as small as 500. The power deteriorates moderately for smaller $N$; false discoveries rise only very slowly. This validates the performance of the test even for small samples.
@@
We turn to measuring feature importance. Table 3 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value. [^11] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities. This is consistent with their treatment in the testing procedure (footnote 9).
@@
The $p$-value distributions show that our approach is able to clearly distinguish between the significant and irrelevant features. The distributions of the significant features $\ell=1,\ldots,12$ are concentrated near 0, far below the level $\alpha=1\%$. This provides substantial evidence in favor of rejecting the null for these features (the rejection probability is equal to 1). The $p$-value distributions of the irrelevant features are concentrated significantly above $\alpha$, implying clear retention of the null for these features. There are no equivocal cases, eliminating the need to actually randomize.
Feature $\ell=9$ is judged to be the most important feature, followed by features $\ell=6,10$ and $12$. The remaining significant features are relatively less important compared with this group of variables. Features $\ell=4,5$ have (almost) identical medians and CIs and thus importance, consistent with their symmetric roles in the conditional mean function (5.1).
To demonstrate the stability of AICO feature importance scores, Figure 4 shows the empirical median (6) for each of the ten training/testing experiments. The AICO importance ranking is stable and robust to noise in the data. The ranking is also robust to alternative choices of the score function $S$ such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request). For comparison, Figure 4 also provides test-implied feature importance scores for CPT, HRT, and LOCO (?, ?, ?), kernel and deep SHAP feature importance (?), permutation feature importance scores (?, ?), and Gaussian and deep knockoff feature importance scores (?, ?). It is important to note that there is no ground truth for feature importance (?); hence we do not focus on the differences in importance rankings across alternative approaches. We do, however, highlight the robustness of the measures based on testing and variable selection procedures. The widely-used SHAP measures deliver inconsistent rankings across implementations (deep vs. kernel). The rankings are also highly unstable across experiments (data sets). That is, they appear to be highly susceptible to noise in the data. This questions the usefulness of SHAP for feature importance.
@@
Finally we consider the computational efficiency of our test vs. the aforementioned alternatives. Figure 5 reports average computation times across 10 experiments. [^12] The compute environment consists of 8GB memory and 4 CPU cores (Intel Xeon CPU E5-2670, 2.60GHz and Silver 4110 CPU, 2.10GHz). In each of these experiments we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features. The time reported includes the time to train any sub or auxiliary models. It does not include the time to train the original neural network (14).
AICO achieves an unprecedented degree of computational efficiency. In terms of full-fledged testing procedures, AICO is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is almost two orders of magnitudes more efficient than Gaussian knockoffs, the fastest screening procedure. It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure. The testing results reported earlier in this section demonstrate that AICOs computational efficiency does not come at the expense of statistical performance.

### v4

@@ 2.2 Multiple features, panel data and missing data
The discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The same framework naturally extends to subsets of $[d]$, enabling the joint testing of multiple features and the exploration of higher-order effects, including feature interactions.
Panel data settings—where collections of units are observed over time—are common in many domains. Appendix B discusses two simple extensions, conditioning and grouping, that adapt the construction of the feature effect to such data structures.
Real-world data sets often contain missing feature values. Appendix C describes how missing or imputed values can be handled when constructing $\Delta_{i}$.
@@ 3 Hypothesis testing
To construct a hypothesis test, we examine the empirical distribution of the feature effects $\Delta_{i}$ across the test samples $i\in I_{2}$, keeping the feature $\ell\in[d]$ fixed. If this distribution has a non-positive center, the feature is deemed insignificant; a large positive center provides evidence of significance. Several classical tests could be applied to assess such central tendencies, including $t$, sign, Wilcoxon, and permutation tests. We propose a simple sign test, which requires minimal assumptions on the data while retaining meaningful asymptotic relative efficiency compared with a $t$-test.
@@
The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$, with unique median $M\in\mathbb{R}$. The associated distribution function $F_{\Delta}$ is continuous. [^5] A sufficient condition for the uniqueness of $M$ is strict monotonicity of $F_{\Delta}$ in a neighborhood of $M$.
Let $B_{n,p}$ be a Binomial-$(n,p)$ variable and let $q_{\lambda}(B_{n,p})=\inf\{x\geq 0:P(B_{n,p}\leq x)\geq\lambda\}$ be its (lower) quantile at $\lambda\in(0,1)$. Let $N=|I_{2}|$ be the size of the test sample, and write $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ for the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
@@
For brevity, we often suppress the dependence of the test statistic $n_{+}$ on $M_{0}$. The procedure rejects $H_{0}$ with probability one if $n_{+}$—the number of feature effects in the test set exceeding $M_{0}$—is greater than the $(1-\alpha)$-quantile of the $\mathrm{Binomial}(N,1/2)$ distribution. If $n_{+}$ lies exactly on the boundary of the rejection region, the test randomizes. Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, this randomization makes the procedure slightly more aggressive: it achieves the maximum possible power while guaranteeing exactness, in the sense that
@@
Algorithm 1 summarizes the implementation of the test. In our empirical applications in Sections 5 and 6, we use the value $M_{0}=0$ in (1).
@@ 4 Measuring feature importance
We seek measures of feature importance that complement the feature $p$-values delivered by our test. A local measure of feature importance is given by the feature effect $\Delta_{i}$, which captures the impact of a feature at the sample level. At the population level, a natural measure of feature importance is the median $M$ of the (conditional) feature-effect distribution $F_{\Delta}$. Under Assumption 3.1, a consistent estimator of $M$ is the empirical median
$\widehat{M}=\mathrm{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$, with approximate equality for even $N$. Hence, (6) is (approximately) as likely to overestimate $M$ as to underestimate it. Moreover, the distribution of $\widehat{M}$ is continuous. The estimator $\widehat{M}$ is unbiased if $F_{\Delta}$ is symmetric about $M$; see Section 4.4 of ?.
@@
We now develop several useful properties of the population-level feature-importance measure defined by the median $M$. Let $Z=(X,Y)\sim\mathbb{P}$, where $\mathbb{P}$ denotes the population law from which the observations $Z_{i}$ are independently drawn. For concreteness, consider a model that estimates the conditional mean, and let $\mu(X)=\mathbb{E}(Y|X)$ with score function $S$ as given. The (population) feature effect is then $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$, where $\underline{X}$ and $\overline{X}$ denote appropriate baseline and comparison features, respectively.
As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers flexibility. Consider the basic choice from Example 2.1,
@@
where, without loss of generality, feature $X^{1}$ is to be tested. Under this specification (10) and given $\mu$ and $S$, the feature effect $\Delta$ is fully determined by the marginal distributions of the features and the joint distribution of the response $Y$ and feature $X^{1}$. Hence, the dependence structure among features does not affect $\Delta$: the importance measure $M$ generated by (10) is robust to feature correlation. This robustness is notable given the well-documented sensitivity of popular importance measures—including SHAP and LOCO—to feature correlation (e.g., ?, ?). Such correlation sensitivity can introduce bias and lead to misleading importance rankings.
Next, we explore the relationship between $M$ and conditional independence. ?, ?, ?, ?, ?, and others develop tests for a conditional-independence hypothesis of feature significance within the model-X knockoff framework of ?. We show that the hypothesis tested by AICO, $M\leq M_{0}=0$, is generally weaker than the alternative conditional-independence hypothesis. Throughout, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
@@
A special case where the converse does hold is the standard linear model
$Y=\mu(X)+\epsilon,\quad\text{where}\quad\mu(X)=\beta\cdot X,\quad\beta=(\beta^{1},\ldots,\beta^{d})\in\mathbb{R}^{d}.$ (11)
Before stating the converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) and assuming symmetry of the distribution of $\Delta$, the median $M$ measuring feature importance is proportional to the square of the corresponding regression coefficient. This is precisely the canonical measure of feature importance in a linear model, with proportionality factor given by the feature’s variance.
@@ 5 Synthetic tasks
This section illustrates the robust performance of our test and feature-importance measure on synthetic “known-truth” regression and classification tasks implemented with neural networks. It benchmarks AICO against existing model-agnostic significance tests and variable-selection procedures, as well as against alternative model-specific and model-agnostic measures of feature importance.
@@ 5.1 Regression
We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ defined by
@@
where the error $\epsilon$ is independent of $X$, and the conditional mean function $\mu$ is given by
@@
The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances, and covariance $0.85$. The features $X_{2},\ldots,X_{5}$ and $X_{7}$ are independent $N(0,1)$ variables, and $X_{8}$ is uniformly distributed on [1,1][-1,1]. The binary feature $X_{9}=1\{X_{2}+Z<0\}$ is generated from an independent $Z\sim N(0,1)$. Feature $X_{10}$ follows a Poisson distribution with mean 3, while $X_{11}$ and $X_{12}$ are independent $t$-distributed variables with five degrees of freedom. The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances, and covariance $0.85$, and $X_{17},X_{18}$, and $X_{19}$ are independent $t$-distributed variables with five degrees of freedom.
We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. We train a single-layer neural network model $f$ to estimate the conditional mean $\mu(X)=\mathbb{E}(Y|X)$, defined as
@@
where $K$ is the number of hidden units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated. Training uses a standard $L_{2}$ loss criterion with $L_{2}$ regularization on the weights $b_{k}$, the Adam optimization method, a batch size of 32, and 500 epochs with early stopping. We iterate across five random initializations. A validation set (25% of $I_{1}$, excluded from training) and a Bayesian search are used to optimize the regularization weight, learning rate, and number of units $K$. The optimal regularization weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$, and the optimal number of units is $K=300$. The resulting model achieves an $R^{2}$ exceeding 99% on both the training and test data.
The testing procedure employs the score function $S(m,y)=-(m-y)^{2}$, consistent with the $L_{2}$ training criterion. We use the basic baseline and comparison feature specifications described in Example 2.1. For testing a feature $\ell\in[d]$, we set
$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i}),$
$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d}),$ (15)
where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ when feature $\ell$ is continuous. For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$, and for the Poisson feature $\ell=10$, we adopt the “alternative mode”
$v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1}),$
where $p^{\ell}(\cdot|I_{1})$ denotes the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. This choice prevents baseline and comparison features from being identical for too many samples, an issue arising with the standard mode. Our specifications ensure that all baseline values lie within the support of $X^{9}$ and $X^{10}$.
@@
We run ten independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report results for confidence levels $\alpha=0.01$ and 0.050.05.777Limited numerical precision can lead to ties, i.e., samples $\Delta_{i}$ whose true values are distinct being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact arises when samples $\Delta_{i}$ with true, nonzero values are rounded to zero. We treat these as nonpositive, excluding them from $n_{+}$—a conservative choice. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. Across all experiments, the test achieves perfect power and produces no false discoveries. Feature correlation (e.g., $\ell=1,6$) and discrete feature distributions (e.g., $\ell=9,10$) do not impair performance. The procedure also correctly identifies the significance of features $\ell=2,4,5$, which influence the response only through interaction terms. Additional experiments confirm robustness to alternative baseline specifications—for example, taking $v_{i}^{\ell}$ as the training mean for the discrete features $\ell=9,10$, or using the median instead of the mean to construct $v_{i}^{\ell}$. We further verify robustness with respect to the score function $S$, including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
@@
For additional perspective, we also implement strong model-agnostic variable-selection procedures, including the Gaussian Model-X Knockoff method of ? and the Deep Model-X Knockoff method of ?. The Gaussian version assumes a multivariate normal feature distribution, whereas the deep version uses a neural network to learn a generative model of the feature distribution. As with CPT and HRT, the construction of such auxiliary models incurs significant additional computational expense, which AICO avoids.
Table 1 reports results for these alternative procedures across ten independent training/testing experiments. CPT, HRT, and LOCO achieve comparable power to AICO but exhibit higher false discovery rates (some of which, but not all, can be mitigated by a Bonferroni correction). The Gaussian Knockoff procedure consistently selects the significant features 1 through 12 in every experiment, though it produces a small number of false discoveries, consistent with its expected false-discovery-rate control. The Deep Knockoff method reduces false discoveries at the cost of lower power and shows modest difficulty identifying features that enter the model (5.1) through the interaction term $2x_{4}x_{5}$.
We examine the performance of the AICO test as a function of the test-set size $N=|I_{2}|$. Table 2 reports results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is shown in Table 1. The test maintains full power for sample sizes as small as $N=500$. For smaller $N$, power declines moderately, while false discoveries increase only marginally. These results confirm that the test performs reliably even in small-sample settings.
@@
We now turn to measuring feature importance. Table 3 presents a feature-importance ranking based on the empirical median (6) for one randomly selected training/testing experiment out of ten. It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, along with the standard (non-randomized) two-sided interval (9) and their corresponding coverage probabilities. In addition, Table 3 provides the distribution of the randomized $p$-values. [^9] Ties among samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing confidence intervals and coverage probabilities. This treatment is consistent with the handling of ties in the testing procedure (footnote 7).
@@
The $p$-value distributions show that our approach clearly separates significant from irrelevant features. For the significant features $\ell=1,\ldots,12$, the distributions are concentrated near zero, far below the level $\alpha=1\%$. This provides strong evidence in favor of rejecting the null hypothesis for these features, with rejection probability equal to one. In contrast, the $p$-value distributions of the irrelevant features are concentrated well above $\alpha$, indicating clear retention of the null. No equivocal cases are observed, eliminating the need for randomization.
Feature $\ell=9$ emerges as the most important variable, followed by features $\ell=6$, $10$, and $12$. The remaining significant features are relatively less influential compared with this group. Features $\ell=4$ and $\ell=5$ exhibit (almost) identical medians and confidence intervals, consistent with their symmetric roles in the conditional mean function (5.1).
To demonstrate the stability of AICO feature-importance scores, Figure 4 reports the empirical median (6) for each of ten independent training/testing experiments. The resulting AICO rankings are stable and robust to noise in the data, and remain consistent under alternative choices of the score function $S$, such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request). For comparison, Figure 4 also displays test-implied feature-importance scores for CPT, HRT, and LOCO (?, ?, ?); kernel and deep SHAP (?); permutation feature-importance scores (?, ?); and Gaussian and deep knockoff importance scores (?, ?).
There is, of course, no ground truth for feature importance (?); we therefore focus on the stability of the rankings rather than their absolute differences across methods. The measures based on formal testing and variable-selection procedures, including AICO, exhibit strong robustness across experiments. In contrast, SHAP delivers inconsistent rankings between its deep and kernel implementations and shows substantial instability across data sets, indicating high sensitivity to noise. These findings raise concerns about the reliability of SHAP-based feature-importance measures.
@@
Finally, we compare the computational efficiency of our test with that of the alternative methods discussed above. Figure 5 reports average computation times across ten experiments. [^10] The compute environment consists of 8GB memory and four CPU cores (Intel Xeon CPU E5-2670, 2.60GHz, and Silver 4110 CPU, 2.10GHz). In each experiment we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features, or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features. The reported time includes any training of submodels or auxiliary models, but excludes the time required to train the original neural network (14).
AICO achieves an unprecedented degree of computational efficiency. Among full-fledged testing procedures, it is more than three orders of magnitude faster than LOCO, the next most efficient test. It is nearly two orders of magnitude faster than Gaussian knockoffs, the fastest screening procedure, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method. As shown earlier in this section, AICO’s exceptional computational efficiency does not come at the expense of statistical performance.